# Nearest Neighbour Score Estimators for Diffusion Generative Models

### v1

@@ 1 Introduction
Diffusion models (Sohl-Dickstein et al., 2015; Ho et al., 2020) have recently emerged as a powerful class of generative models. They have seen broad adoption across a variety of tasks such as image generation (Rombach et al., 2022), video generation (Harvey et al., 2022), and 3D object synthesis (Poole et al., 2023). Although these models achieve state of the art performance, their sampling procedure requires integration of the probability flow ODE (PF-ODE) or diffusion SDE (Song et al., 2021). This procedure hampers sampling speed as each integration step requires a neural network evaluation. Typically, many integration steps are required per sample (Ho et al., 2020; Karras et al., 2022).
@@
Motivated by this shortcoming, an array of methods have been proposed which learn few-step approximations of the PF-ODE solution. Amongst these, some attempt to distill a pre-trained diffusion network such as progressive distillation (Salimans & Ho, 2022), consistency distillation (Song et al., 2023) or adversarial diffusion distillation (Sauer et al., 2023). Alternatively, consistency training (Song et al., 2023; Song & Dhariwal, 2023) proposes a method to train consistency models without a pre-trained diffusion model. We refer to the general class of models including diffusion models and related few-step methods as diffusion generative models.
@@
As diffusion generative models are grounded in diffusion processes, they utilize the score function. This critical quantity can be estimated by network approximation or Monte Carlo methods. Diffusion and consistency training utilize a one-sample Monte Carlo estimator of the score (excepting (Xu et al., 2023)) in their training procedure. Alternatively, both diffusion sampling and diffusion distillation methods rely upon neural network estimators. However, each estimator has drawbacks. The Monte Carlo estimator has high variance (Xu et al., 2023) while the neural network estimator is imperfect, resulting in bias (Karras et al., 2022).
@@
Our work puts forward a new method for score function estimation with lower variance than the one-sample estimator and less bias than the neural network approximation. Utilizing self-normalized importance sampling (Hesterberg, 1995), we estimate the score through a weighted average over a batch of training set examples We draw these elements from a proposal which prioritizes examples which are the most likely to have generated the current noisy value (Figure 1). We show that because diffusion processes are Gaussian, these examples are equivalent to the $\ell_{2}$ nearest neighbours to the noisy data. With this finding, we can rapidly identify and sample important examples using a fast KNN search (Johnson et al., 2019). We analyze our method and derive bounds on the variance of our estimator.
@@
Empirically, we measure our performance in three settings. First, we compare our score estimate against the true score on CIFAR-10 (Krizhevsky et al., 2009). We find that our method has near-zero variance and bias – substantially outperforming both STF (Xu et al., 2023) and EDM (Karras et al., 2022). Applying our method to consistency models, we find that replacing one-sample estimators with our method improves consistency training – resulting in faster convergence and higher sample quality. Finally, we show that our method can replace neural networks in PF-ODE integration, opening interesting future research avenues.
@@ 2 Background
Since the density $p(\mathbf{x})$ is usually unknown, it is approximated by $p_{data}(\mathbf{x})$, a finite mixture of Dirac measures $p_{data}\left(\mathbf{x}\right)=\frac{1}{N}\sum_{\mathbf{x}^{(i)}\in\mathcal{D}}\delta\left(\mathbf{x}-\mathbf{x}^{(i)}\right)$ where $\mathcal{D}=\left\{\mathbf{x}^{(1)},\ldots,\mathbf{x}^{(N)}\right\}$ is a dataset of samples drawn from the true data density. With this approximation, the time-varying marginals $p_{t}(\mathbf{z})$ and posteriors $p_{t}(\mathbf{x}^{(i)}|\mathbf{z})$ take the forms
$p_{t}(\mathbf{z})=\sum_{i=1}^{N}p_{t}(\mathbf{z}|\mathbf{x}^{(i)})p_{data}(\mathbf{x}^{(i)})=\sum_{i=1}^{N}\frac{p_{t}(\mathbf{z}|\mathbf{x}^{(i)})}{N}$ (1)
$p_{t}(\mathbf{x}^{(i)}|\mathbf{z})=\frac{p_{t}(\mathbf{z}|\mathbf{x}^{(i)})p_{data}\left(\mathbf{x}^{(i)}\right)}{p_{t}(\mathbf{z})}=\frac{p_{t}(\mathbf{z}|\mathbf{x}^{(i)})}{N\cdot p_{t}(\mathbf{z})}$ (2)
Under the finite data approximation, each posterior distribution $p_{t}(\mathbf{x}^{(i)}|\mathbf{z})$ is categorical over the elements of $\mathcal{D}$, ie. $\sum_{i=1}^{N}p(\mathbf{x}^{(i)}|\mathbf{z})=1$. In addition, we refer to $p_{t}(\mathbf{z}|\mathbf{x}^{(i)})=\mathcal{N}\left(\mathbf{z}|\mathbf{x},\sigma(t)^{2}\mathbf{I}\right)$ as the forward likelihood.
Diffusion processes can be described using stochastic or ordinary differential equations (Song et al., 2021). In this work, we focus on the probability flow ODE (PF-ODE) formulation, introduced by Song et al. (2021). Following the parameterization of EDM (Karras et al., 2022), we select $\sigma(t)=t$ although other choices are possible, as discussed in Appendix B. The PF-ODE for this parameterization is
@@ 2.1 Diffusion Models
Diffusion models are a class of generative models which can generate samples through numerical integration of Equation 3 from $T$ to $t_{\mathrm{min}}$ with initial value $\mathbf{z}\sim\pi(\mathbf{\mathbf{z}})$.
@@
The minimizer of Equation 6 is $D^{\star}_{\theta}\left(\mathbf{z},t\right)=\mathbb{E}\left[\mathbf{x}|\mathbf{z},t\right]$ (Karras et al., 2022). However, during SGD training, per-batch loss is calculated against $(\mathbf{x}^{(i)},\mathbf{z})$ tuples drawn from the joint distribution instead of against the true posterior mean $\mathbb{E}[\mathbf{x}|\mathbf{z},t]$. Minimizing the loss with individual $\mathbf{x}^{(i)}$ converges to $\mathbb{E}\left[\mathbf{x}|\mathbf{z},t\right]$ because the joint distribution from which the tuples are drawn $p_{t}(\mathbf{z},\mathbf{x}^{(i)})$ is proportional to $p_{t}(\mathbf{x}^{(i)}|\mathbf{z})$. Due to this relation, each $\mathbf{x}^{(i)}$ can be seen as a single-sample Monte Carlo estimator of the posterior mean (Xu et al., 2023).
@@ 2.2 Consistency Models
Equation 3 defines trajectories which map between any marginal distribution $p_{t}(\mathbf{z})$ and $p_{t_{\mathrm{min}}}(\mathbf{z})\approx p_{data}\left(\mathbf{x}\right)$, where $t_{\mathrm{min}}$ is used for numerical stability. These mappings $\bm{f}:(\mathbf{z},t)\rightarrow\mathbf{x}$ are known as consistency functions because the value $\bm{f}(\mathbf{z},t)=\mathbf{x}^{\star}$ is consistent for all $(\mathbf{z},t)$ on a common PF-ODE trajectory. Consistency models (Song et al., 2023) exploit this self-consistency property to learn $\bm{f}_{\bm{\theta}}(\mathbf{z},t)$, a neural network parameterized approximation to the true consistency function $\bm{f}$. In practice, consistency models are trained by optimizing the consistency matching objective $\mathcal{L}_{CM}=\mathbb{E}_{t_{i}}\left[\lambda(t_{i})\mathcal{L}_{t_{i}}\right]$ where
@@
The loss aims to ensure consistency between the student network $\bm{f}_{\theta}$ and the teacher network $\bm{f}_{\theta^{-}}$, evaluated at points $(\mathbf{z},t_{i})$ and $(\hat{\mathbf{z}},t_{i-1})$ which lie on the same PF-ODE trajectory. To obtain $\hat{\mathbf{z}}$, consistency models perform a one step integration of Equation 3 from from $t_{i}$ to $t_{i-1}$, setting $\hat{\mathbf{z}}=\texttt{solver}\left(\mathbf{z},t_{i},t_{i-1},\nabla_{z}\log p_{t_{i}}(\mathbf{z})\right)$7. The score function is therefore critical to performing this integration and ensuring $\mathbf{z}$ and $\hat{\mathbf{z}}$ lie on the same PF-ODE path.
Since $\nabla_{z}\log p_{t_{i}}(\mathbf{z})$ is infeasible to calculate exactly outside the small data regime, consistency models are trained with score estimators. Song et al. (2023) propose training consistency models with two such estimators. Consistency distillation utilizes a pre-trained neural network approximator $\nabla_{z}\log p_{t_{i}}(\mathbf{z})\approx s_{\theta}(\mathbf{z},t_{i})$. To train consistency models without a pre-trained score estimator, consistency training utilizes the estimator $\nabla_{z}\log p_{t_{i}}(\mathbf{z})\approx\frac{\mathbf{x}-\mathbf{z}}{t^{2}}$ with $\mathbf{x}$ and $\mathbf{z}$ sampled from the joint $p_{t}(\mathbf{z},\mathbf{x}).$ Like diffusion training, the full posterior expectation is replaced in this case with a single-sample Monte Carlo estimate.
@@ 3.1 SNIS Score Estimation
Although the score is necessary for both training and sampling from diffusion generative models, there are only two widely used estimators of the score function - learned networks or single-sample Monte Carlo estimators. Since the latter is required to produce the former, we consider methods to reduce variance in Monte Carlo score estimation.
@@
To increase $n$ and thereby decrease estimator variance, we appeal to importance sampling (IS)(Kahn & Marshall, 1953) taking inspiration from Xu et al. (2023). Instead of sampling from $p_{t}(\mathbf{x}^{(i)}|\mathbf{z})$, we draw samples from a proposal distribution $q_{t}(\mathbf{x}^{(i)}|\mathbf{z})$. We can formulate a new Monte Carlo estimator using a batch of $n$ samples drawn from the proposal $S_{q}=\left\{\mathbf{x}_{1},\ldots,\mathbf{x}_{n}\right\}\sim q_{t}(\mathbf{x}^{(i)}|\mathbf{z})$ as
$\hat{\mu}_{\text{IS}}=\frac{1}{n}\sum_{\mathbf{x}^{(i)}\in S_{q}}\frac{p_{t}\left(\mathbf{z}|\mathbf{x}^{(i)}\right)p_{data}\left(\mathbf{x}^{(i)}\right)}{q_{t}(\mathbf{x}^{(i)}|\mathbf{z})p_{t}\left(\mathbf{z}\right)}\mathbf{x}^{(i)}.$ (10)
Equation 10 no longer requires samples from the posterior and with the finite data approximation, we know $p_{data}(\mathbf{x}^{(i)})=1/N$. However, finding $p_{t}(\mathbf{z})$ still requires a sum over $\mathcal{D}$. We resolve this issue with self-normalized importance sampling (Hesterberg, 1995). By also estimating Equation 1 with an IS estimator using the same $S_{q}$, we arrive at a self-normalized importance sampling (SNIS) estimator
$\hat{\mu}_{\text{SNIS}}=\frac{\sum_{\mathbf{x}^{(i)}\in S_{q}}w_{i}\mathbf{x}^{(i)}}{\sum_{\mathbf{x}^{(j)}\in S_{q}}w_{j}},\ w_{i}=\frac{p_{t}(\mathbf{z}|\mathbf{x}^{(i)})}{q_{t}(\mathbf{x}^{(i)}|\mathbf{z})}.$ (11)
The constant factor $p_{data}(\mathbf{x}^{(i)})$ cancels because it is present in the numerator of Equation 10 and in Equation 1.
@@ 3.2 Nearest Neighbour Proposals
With a defined a framework for reducing score estimator variance, we must next determine a suitable proposal distribution $q_{t}(\mathbf{x}^{(i)}|\mathbf{z})$. To identify a good proposal, we start by analyzing the variance of Equation 11 (Owen, 2013). The diagonal of the SNIS score estimator covariance is
$\begin{split}\text{Diag}&\left(\text{Cov}\left(\frac{\hat{\mu}_{\text{SNIS}}-\mathbf{z}}{t^{2}}\right)\right)\\ &=\frac{1}{nt^{4}}\sum_{i=1}^{N}\frac{p_{t}\left(\mathbf{x}^{(i)}|\mathbf{z}\right)^{2}}{q_{t}\left(\mathbf{x}^{(i)}|\mathbf{z}\right)}\left(\mathbf{x}^{(i)}-\mu\right)^{\circ 2}\end{split}$ (12)
Where ${}^{\circ 2}$ indicates the Hadamard power operator.
Examining Equation 12, we can infer that an appropriate proposal distribution $q_{t}(\mathbf{x}^{(i)}|\mathbf{z})$ is important for estimating $\mathbb{E}\left[\mathbf{x}|\mathbf{z},t\right]$ with low per-dimension variance. Specifically, high importance ratios $p_{t}(\mathbf{x}^{(i)}|\mathbf{z})/q_{t}(\mathbf{x}^{(i)}|\mathbf{z})$ will increase variance if $(\mathbf{x}^{(i)}-\mu)^{\circ 2}$ is non-zero.
@@
Although we cannot make immediate use of the univariate optimal proposal, we note that $q^{\star}$ is proportional to the target distribution $p(x)$. We therefore hypothesize that $q_{t}(\mathbf{x}^{(i)}|\mathbf{z})=p_{t}(\mathbf{x}^{(i)}|\mathbf{z})$ is a reasonable choice of proposal distribution. Because $p_{t}(\mathbf{x}^{(i)}|\mathbf{z})$ varies with $t$ and $\mathbf{z}$, we also vary $q_{t}(\mathbf{x}^{(i)}|\mathbf{z})$ with these variables.
Determining the posterior exactly requires summing over $\mathcal{D}$ to compute $p_{t}(\mathbf{z})$. Instead, we structure $q_{t}(\mathbf{x}^{(i)}|\mathbf{z})$ as a $k$ nearest neighbour truncated approximation of $p_{t}(\mathbf{x}^{(i)}|\mathbf{z})$. Let $\mathcal{K}(\mathbf{z},t)=(\mathbf{x}_{1},\ldots\mathbf{x}_{k}|\mathbf{x}_{i}\in\mathcal{D})\ s.t.\ p_{t}(\mathbf{x}_{1}|\mathbf{z})\geq\ldots\geq p_{t}(\mathbf{x}_{k}|\mathbf{z})$ be the ordered set of the $k$ most probable dataset elements under the posterior distribution $p_{t}(\mathbf{x}^{(i)}|\mathbf{z})$. The subscript of each $\mathbf{x}_{j}\in\mathcal{K}(\mathbf{z},t)$ indicates its ordering, where $\mathbf{x}_{k}$ is the element of $\mathcal{K}$ with the lowest probability. For brevity, we refer to $\mathcal{K}(\mathbf{z},t)$ as $\mathcal{K}$ hereafter and use it to define our nearest neighbours proposal
$\mathcal{Z}_{q}\cdot q_{t}(\mathbf{x}^{(i)}|\ \mathbf{z})=\begin{cases}p_{t}\left(\mathbf{z}|\mathbf{x}^{(i)}\right)&\mathbf{x}^{(i)}\in\mathcal{K}\\ p_{t}\left(\mathbf{z}|\mathbf{x}_{k}\right)&\mathbf{x}^{(i)}\notin\mathcal{K}\end{cases}.$ (13)
There are several beneficial properties of our nearest neighbours proposal. Since $p_{t}(\mathbf{z}|\mathbf{x})=\mathcal{N}(\mathbf{z}|\mathbf{x},t^{2}\mathbf{I})$ is a Gaussian, the forward likelihood of each $\mathbf{x}^{(i)}$ can be computed easily. Additionally, since the size of $\mathcal{D}$ is known, the normalizing constant $\mathcal{Z}_{q}=\sum_{i=1}^{k}p_{t}(\mathbf{z}|\mathbf{x}_{i})+(N-k)p_{t}(\mathbf{z}|\mathbf{x}_{k})$ is found easily. Finally, the shape of $q_{t}(\mathbf{x}|\mathbf{z})$ matches the shape of $p_{t}(\mathbf{x}|\mathbf{z})$, as demonstrated in Figure 1.
@@
  Input: $\mathbf{z}$, $t$, dataset $\mathcal{D}$, index $\mathcal{I}$
  $\mathcal{K},d\leftarrow\texttt{search}(\mathcal{I},\mathbf{z})$
  $q_{t}(\mathbf{x}^{(i)}|\mathbf{z})\leftarrow\text{exp}(\frac{-d_{i}^{2}}{2t^{2}})\ \forall\mathbf{x}^{(i)}\in\mathcal{K}$Equation 13}
  $q_{t}(\mathbf{x}^{(i)}|\mathbf{z})\leftarrow\text{exp}(\frac{-d_{k}^{2}}{2t^{2}})\ \forall\mathbf{x}^{(i)}\notin\mathcal{K}$
  $q_{t}(\mathbf{x}^{(i)}|\mathbf{z})\leftarrow\frac{q(\mathbf{x}^{(i)}|\mathbf{z})}{\sum_{j=1}^{|\mathcal{K}|}q(\mathbf{x}^{(i)}|\mathbf{z})}$
@@
  $\hat{\mu}_{\text{KNN}}\leftarrow\sum_{i=1}^{n}\bar{w}_{i}\mathbf{x}_{i}$Equation 11}
@@
For $\mathbf{x}^{(i)}\in\mathcal{K}$, the proposal is proportional to the posterior with $\mathcal{Z}_{q}q_{t}(\mathbf{x}|\mathbf{z})=p_{t}(\mathbf{x})p_{t}(\mathbf{x}|\mathbf{z})$. In general, $p_{t}(\mathbf{z})\leq\mathcal{Z}_{q}$ with equality in two cases. When $p_{t}(\mathbf{z}|\mathbf{x}_{k})\approx 0$ or $k=N$, the posterior is fully concentrated on $\mathcal{K}$ and $p_{t}(\mathbf{z})=\sum_{i=1}^{N}p_{t}(\mathbf{z}|\mathbf{x}^{(i)})=\sum_{\mathbf{x}^{(i)}\in\mathcal{K}}p_{t}(\mathbf{z}|\mathbf{x}^{(i)})$. When $p_{t}(\mathbf{z}|\mathbf{x}_{k})>0$ and $k<N$, the likelihood $p_{t}(\mathbf{z}|\mathbf{x}^{(i)})$ upper bounds the likelihood for all $\mathbf{x}^{(i)}\notin\mathcal{K}$. In this case, we underweight the probability for $\mathbf{x}^{(i)}\in\mathcal{K}$, with more mass placed in the tails of the proposal distribution.

### v2

@@ 1 Introduction
Diffusion models (Sohl-Dickstein et al., 2015; Ho et al., 2020; Song et al., 2021) have emerged as a powerful class of generative models. They have seen broad adoption across a variety of tasks such as image generation (Rombach et al., 2022), video generation (Harvey et al., 2022), and 3D object synthesis (Poole et al., 2023). Although these models achieve state of the art performance, their sampling procedure requires integration of the probability flow ODE (PF-ODE) or diffusion SDE (Song et al., 2021). This procedure hampers sampling speed as each integration step requires a neural network evaluation. Typically, many integration steps are required per sample (Ho et al., 2020; Karras et al., 2022).
@@
Motivated by this shortcoming, an array of methods have been proposed which learn few-step approximations of the PF-ODE solution. Amongst these, some attempt to distill a pre-trained diffusion network such as progressive distillation (Salimans & Ho, 2022), consistency distillation (Song et al., 2023) or adversarial diffusion distillation (Sauer et al., 2023). Alternatively, consistency training (Song et al., 2023; Song & Dhariwal, 2024) proposes a method to train consistency models without a pre-trained diffusion model. We refer to the general class of models including diffusion models and related few-step methods as diffusion generative models.
@@
As diffusion generative models are grounded in diffusion processes, they utilize the score function. This critical quantity can be estimated by network approximation or Monte Carlo methods. Diffusion and consistency training utilize a one-sample Monte Carlo estimator of the score in their training procedure, (excepting (Xu et al., 2023)). Alternatively, both diffusion sampling and diffusion distillation methods rely upon neural network estimators. However, each estimator has drawbacks. The single-sample Monte Carlo estimator has high variance (Xu et al., 2023) while learned neural network approximators are imperfect, leading to bias (Karras et al., 2022).
@@
Our work puts forward a new method for score function estimation with lower variance than the one-sample estimator and less bias than neural network approximation. Utilizing self-normalized importance sampling (Hesterberg, 1995), we estimate the score through a weighted average over a batch of training set examples. We draw these elements from a proposal which prioritizes examples which are the most likely to have generated the current noisy value (Figure 1). We show that because diffusion processes are Gaussian, these examples are equivalent to the $\ell_{2}$ nearest neighbours to the noisy data. With this finding, we can rapidly identify and sample important examples using a fast KNN search (Johnson et al., 2019). We analyze our method and derive bounds on the variance of our estimator.
@@
Empirically, we measure our performance in three settings. First, we compare our score estimate against the analytic score on CIFAR-10 (Krizhevsky et al., 2009). We find that our method has near-zero variance and bias – substantially outperforming both STF (Xu et al., 2023) and EDM (Karras et al., 2022). Applying our method to consistency models, we find that replacing one-sample estimators with our method improves consistency training – resulting in faster convergence and higher sample quality. Finally, we show that our method can replace neural networks in PF-ODE integration, opening interesting future research avenues. We release our code [^1] https://github.com/plai-group/knn-score to encourage use of our estimator by the community.
@@ 2 Background
Since the density $p(\mathbf{x})$ is usually unknown, it is approximated by $p_{\text{data}}(\mathbf{x})$, a finite mixture of Dirac measures $p_{data}\left(\mathbf{x}\right)=\frac{1}{N}\sum_{\mathbf{x}^{(i)}\in\mathcal{D}}\delta\left(\mathbf{x}-\mathbf{x}^{(i)}\right)$ where $\mathcal{D}=\left\{\mathbf{x}^{(1)},\ldots,\mathbf{x}^{(N)}\right\}$ is a dataset of samples drawn from the true data density. With this approximation, the time-varying marginals $p_{t}(\mathbf{z})$ and posteriors $p_{t}(\mathbf{x}^{(i)}|\mathbf{z})$ take the forms
$p_{t}(\mathbf{z})=\sum_{i=1}^{N}p_{t}(\mathbf{z}|\mathbf{x}^{(i)})p_{\text{data}}(\mathbf{x}^{(i)})=\sum_{i=1}^{N}\frac{p_{t}(\mathbf{z}|\mathbf{x}^{(i)})}{N}$ (1)
$p_{t}(\mathbf{x}^{(i)}|\mathbf{z})=\frac{p_{t}(\mathbf{z}|\mathbf{x}^{(i)})p_{\text{data}}\left(\mathbf{x}^{(i)}\right)}{p_{t}(\mathbf{z})}=\frac{p_{t}(\mathbf{z}|\mathbf{x}^{(i)})}{N\cdot p_{t}(\mathbf{z})}$ (2)
Under the finite data approximation, each posterior distribution $p_{t}(\mathbf{x}^{(i)}|\mathbf{z})$ is categorical over the elements of $\mathcal{D}$, ie. $\sum_{i=1}^{N}p(\mathbf{x}^{(i)}|\mathbf{z})=1$. In addition, we refer to $p_{t}(\mathbf{z}|\mathbf{x}^{(i)})=\mathcal{N}\left(\mathbf{z};\mathbf{x},\sigma(t)^{2}\mathbf{I}\right)$ as the forward likelihood.
Diffusion processes can be described using stochastic or ordinary differential equations (Song et al., 2021). In this work, we focus on the probability flow ODE (PF-ODE) formulation, introduced by Song et al. (2021). Following the parameterization of EDM (Karras et al., 2022), we select $\sigma(t)=t$. However, other choices are possible, as discussed in Appendix B. With our chosen diffusion schedule, the PF-ODE is given by
@@ 2.1 Diffusion Models
Diffusion models are a class of generative models which can generate samples through numerical integration of Equation 3 from $T$ to $t_{\mathrm{min}}\approx 0$ with initial value $\mathbf{z}\sim\pi(\mathbf{\mathbf{z}})$.
@@
The minimizer of Equation 6 is $D^{\star}_{\theta}\left(\mathbf{z},t\right)=\mathbb{E}\left[\mathbf{x}|\mathbf{z},t\right]$ (Karras et al., 2022). However, during stochastic gradient descent optimization, per-batch loss is calculated using $(\mathbf{x}^{(i)},\mathbf{z})$ tuples drawn from the joint distribution instead of against the true posterior mean $\mathbb{E}[\mathbf{x}|\mathbf{z},t]$. Minimizing the loss with individual $\mathbf{x}^{(i)}$ converges to $\mathbb{E}\left[\mathbf{x}|\mathbf{z},t\right]$ because the joint distribution from which the tuples are drawn $p_{t}(\mathbf{z},\mathbf{x}^{(i)})$ is proportional to $p_{t}(\mathbf{x}^{(i)}|\mathbf{z})$. Due to this relation, each $\mathbf{x}^{(i)}$ can be seen as a single-sample Monte Carlo estimator of the posterior mean (Xu et al., 2023).
@@ 2.2 Consistency Models
Equation 3 defines trajectories which map between any marginal distribution $p_{t}(\mathbf{z})$ and $p_{t_{\mathrm{min}}}(\mathbf{z})\approx p_{data}\left(\mathbf{x}\right)$, where $t_{\mathrm{min}}\approx 0$ is used for numerical stability. These mappings $\bm{f}:(\mathbf{z},t)\rightarrow\mathbf{x}$ are known as consistency functions because the value $\bm{f}(\mathbf{z},t)=\mathbf{x}^{\star}$ is consistent for all $(\mathbf{z},t)$ on a common PF-ODE trajectory. Consistency models (Song et al., 2023) exploit this self-consistency property to learn $\bm{f}_{\bm{\theta}}(\mathbf{z},t)$, a neural network parameterized approximation to the true consistency function $\bm{f}$. In practice, consistency models are trained by optimizing the consistency matching objective $\mathcal{L}_{CM}=\mathbb{E}_{t_{i}}\left[\lambda(t_{i})\mathcal{L}_{t_{i}}\right]$ where
@@
The loss aims to ensure consistency between the student network $\bm{f}_{\theta}$ and the teacher network $\bm{f}_{\theta^{-}}$, evaluated at points $(\mathbf{z},t_{i})$ and $(\hat{\mathbf{z}},t_{i-1})$ which lie on the same PF-ODE trajectory. To obtain $\hat{\mathbf{z}}$, consistency model training performs a one step integration of Equation 3 from from $t_{i}$ to $t_{i-1}$, setting $\hat{\mathbf{z}}=\texttt{solver}\left(\mathbf{z},t_{i},t_{i-1},\nabla_{\mathbf{z}}\log p_{t_{i}}(\mathbf{z})\right)$. An accurate score function estimate is critical to performing this integration, ensuring $\mathbf{z}$ and $\hat{\mathbf{z}}$ lie on the same PF-ODE path.
Since $\nabla_{\mathbf{z}}\log p_{t_{i}}(\mathbf{z})$ is infeasible to calculate exactly outside the small data regime, consistency models are trained with score estimators. Song et al. (2023) propose training consistency models with two such estimators. Consistency distillation utilizes a neural network approximator $\nabla_{\mathbf{z}}\log p_{t_{i}}(\mathbf{z})\approx s_{\theta}(\mathbf{z},t_{i})$ from a pre-trained diffusion “teacher” model. To train consistency models without a pre-trained score estimator, consistency training utilizes the estimator $\nabla_{\mathbf{z}}\log p_{t_{i}}(\mathbf{z})\approx\frac{\mathbf{x}^{(i)}-\mathbf{z}}{t^{2}}$ with $\mathbf{x}^{(i)}$ and $\mathbf{z}$ sampled from $p_{t}(\mathbf{z},\mathbf{x}^{(i)}).$ Like diffusion training, a single-sample Monte Carlo estimate is used in place of the marginal score.
@@ 3.1 SNIS Score Estimation
Although the score is necessary for both training and sampling from diffusion generative models, there are only two widely used estimators of the score function – learned networks or single-sample Monte Carlo estimators. Since the latter is required to produce the former, we consider methods to reduce variance in Monte Carlo score estimation.
@@
To increase $n$ and thereby decrease estimator variance, we appeal to importance sampling (IS)(Kahn & Marshall, 1953) taking inspiration from Xu et al. (2023). Instead of sampling from $p_{t}(\mathbf{x}^{(i)}|\mathbf{z})$, we draw samples from a proposal distribution $q(\mathbf{x}^{(i)})$. We can formulate a new Monte Carlo estimator using a batch of $n$ samples drawn from the proposal $S_{q}=\left\{\mathbf{x}_{1},\ldots,\mathbf{x}_{n}\right\}\sim q(\mathbf{x}^{(i)})$ as
$\hat{\mu}_{\text{IS}}=\frac{1}{n}\sum_{\mathbf{x}^{(i)}\in S_{q}}\frac{p_{t}\left(\mathbf{z}|\mathbf{x}^{(i)}\right)p_{\text{data}}\left(\mathbf{x}^{(i)}\right)}{p_{t}\left(\mathbf{z}\right)q(\mathbf{x}^{(i)})}\mathbf{x}^{(i)}.$ (10)
Equation 10 no longer requires samples from the posterior and with the finite data approximation, we know $p_{\text{data}}(\mathbf{x}^{(i)})=1/N$. However, finding $p_{t}(\mathbf{z})$ still requires a sum over $\mathcal{D}$. We resolve this issue with self-normalized importance sampling (Hesterberg, 1995). By also estimating $p_{t}(\mathbf{z})$ with an IS estimator of Equation 1 using the same $S_{q}$, we arrive at a self-normalized importance sampling (SNIS) estimator
$\hat{\mu}_{\text{SNIS}}=\frac{\sum_{\mathbf{x}^{(i)}\in S_{q}}w_{i}\mathbf{x}^{(i)}}{\sum_{\mathbf{x}^{(j)}\in S_{q}}w_{j}},\ w_{i}=\frac{p_{t}(\mathbf{z}|\mathbf{x}^{(i)})}{q(\mathbf{x}^{(i)})}.$ (11)
The constant factor $p_{\text{data}}(\mathbf{x}^{(i)})$ cancels because it is present in the numerator of Equation 10 and in Equation 1.
@@ 3.2 Nearest Neighbour Proposals
With a defined a framework for reducing score estimator variance, we must next determine a suitable proposal distribution $q(\mathbf{x}^{(i)})$. To identify a good proposal, we start by analyzing the variance of Equation 11 (Owen, 2013). The diagonal of the SNIS score estimator covariance is
$\begin{split}\text{Diag}&\left(\text{Cov}\left(\frac{\hat{\mu}_{\text{SNIS}}-\mathbf{z}}{t^{2}}\right)\right)\\ &=\frac{1}{nt^{4}}\sum_{i=1}^{N}\frac{p_{t}\left(\mathbf{x}^{(i)}|\mathbf{z}\right)^{2}}{q\left(\mathbf{x}^{(i)}\right)}\left(\mathbf{x}^{(i)}-\mu\right)^{\circ 2}\end{split}$ (12)
Where ∘2 indicates the Hadamard power operator.
Examining Equation 12, we can infer that an appropriate proposal distribution $q(\mathbf{x}^{(i)})$ is important for estimating $\mathbb{E}\left[\mathbf{x}|\mathbf{z},t\right]$ with low per-dimension variance. Specifically, high importance ratios $p_{t}(\mathbf{x}^{(i)}|\mathbf{z})/q(\mathbf{x}^{(i)})$ will increase variance if $(\mathbf{x}^{(i)}-\mu)^{\circ 2}$ is non-zero.
@@
Although we cannot make immediate use of the univariate optimal proposal, we note that $q^{\star}$ is proportional to the target distribution $p(x)$. We therefore hypothesize that a suitable choice for our proposal is to match the true posterior $p_{t}(\mathbf{x}^{(i)}|\mathbf{z})$ as closely as possible. Because $p_{t}(\mathbf{x}^{(i)}|\mathbf{z})$ varies with $t$ and $\mathbf{z}$, we propose also varying our proposal with these variables, denoting our proposal as $q_{t}(\mathbf{x}^{(i)}|\mathbf{z})$.
Determining the posterior exactly requires summing over $\mathcal{D}$ to compute $p_{t}(\mathbf{z})$. To more efficiently construct our proposal, we structure $q_{t}(\mathbf{x}^{(i)}|\mathbf{z})$ to approximate $p_{t}(\mathbf{x}^{(i)}|\mathbf{z})$ by using only $k$ most probable elements of $p_{t}(\mathbf{x}^{(i)}|\mathbf{z})$. Let $\mathcal{K}(\mathbf{z},t)=(\mathbf{x}_{1},\ldots\mathbf{x}_{k}|\mathbf{x}_{i}\in\mathcal{D})\ s.t.\ p_{t}(\mathbf{x}_{1}|\mathbf{z})\geq\ldots\geq p_{t}(\mathbf{x}_{k}|\mathbf{z})$ be the ordered set of the $k$ most probable dataset elements under the posterior distribution $p_{t}(\mathbf{x}^{(i)}|\mathbf{z})$. The subscript of each $\mathbf{x}_{j}\in\mathcal{K}(\mathbf{z},t)$ indicates its ordering, where $\mathbf{x}_{k}$ is the element of $\mathcal{K}(\mathbf{z},t)$ with the lowest probability. For brevity, we refer to $\mathcal{K}(\mathbf{z},t)$ as $\mathcal{K}$ hereafter and use it to define our nearest neighbours proposal
$q_{t}(\mathbf{x}^{(i)}|\ \mathbf{z})=\frac{1}{\mathcal{Z}_{q}}\begin{cases}p_{t}\left(\mathbf{z}|\mathbf{x}^{(i)}\right)&\mathbf{x}^{(i)}\in\mathcal{K}\\ p_{t}\left(\mathbf{z}|\mathbf{x}_{k}\right)&\mathbf{x}^{(i)}\notin\mathcal{K}\end{cases}.$ (13)
@@
  Input: $\mathbf{z}$, $t$, $\mathcal{D}$, $\mathcal{I}$,$k$,$n$
  $\mathcal{K},d\leftarrow\texttt{search}(\mathcal{I},\mathbf{z},k)$
  $p_{t}(\mathbf{z}|\mathbf{x}_{i})\leftarrow\text{exp}(\frac{-d_{i}^{2}}{2t^{2}})\ \forall\ \mathbf{x}_{i}\in\mathcal{K}$
  $\mathcal{Z}_{q}\leftarrow\sum_{\mathbf{x}_{i}}p_{t}(\mathbf{z}|\mathbf{x}_{i})+(N-k)p_{t}(\mathbf{z}|\mathbf{x}_{k})$
  $q_{t}(\mathbf{x}^{(i)}|\mathbf{z})\leftarrow\frac{1}{\mathcal{Z}_{q}}\begin{cases}&p_{t}(\mathbf{z}|\mathbf{x}_{i})\ \forall\ \mathbf{x}^{(i)}\in\mathcal{K}\\ &p_{t}(\mathbf{z}|\mathbf{x}_{k})\ \forall\ \mathbf{x}^{(i)}\notin\mathcal{K}\end{cases}$
@@
  $\hat{\mu}_{\text{KNN}}\leftarrow\sum_{i=1}^{n}\bar{w}_{i}\mathbf{x}_{i}$